# The dark side of the forces: assessing non-conservative force models for atomistic machine learning

### 19 Feb 2025 v2

@@ Abstract
The use of machine learning to estimate the energy of a group of atoms, and the forces that drive them to more stable configurations, have revolutionized the fields of computational chemistry and materials discovery. In this domain, rigorous enforcement of symmetry and conservation laws has traditionally been considered essential. For this reason, interatomic forces are usually computed as the derivatives of the potential energy, ensuring energy conservation. Several recent works have questioned this physically constrained approach, suggesting that directly predicting the forces yields a better trade-off between accuracy and computational efficiency – and that energy conservation can be learned during training. This work investigates the applicability of such non-conservative models in microscopic simulations. We identify and demonstrate several fundamental issues, from ill-defined convergence of geometry optimization to instability in various types of molecular dynamics. Contrary to the case of rotational symmetry, energy conservation is hard to learn, monitor, and correct for. The best approach to exploit the acceleration afforded by direct force prediction might be to use it in tandem with a conservative model, reducing – rather than eliminating – the additional cost of backpropagation, but avoiding the pathological behavior associated with non-conservative forces.
Machine Learning, ICML
@@ 1 Introduction
Interatomic potentials model the microscopic interactions between atoms, and determine – directly or by modulating the response to thermal excitations – the stability and reactivity of molecules and materials. Over many decades, interatomic potentials have been used in Monte Carlo (Metropolis et al., 1953) and molecular dynamics (MD) (Andersen, 1980) simulations, geometry optimization, and other techniques, allowing a mechanistic study of the atomic-scale behavior and properties of molecules, materials, and biological systems (Allen & Tildesley, 2017; Tuckerman, 2008). While traditional interatomic potentials are based on simple physically inspired functional forms, the last decade has seen machine-learned interatomic potentials (MLIPs) obtaining remarkable accuracies at a high level of computational efficiency, most often by learning reference energies and forces from much slower first-principle quantum mechanical calculations (Behler, 2021; Unke et al., 2021).
@@ 2.1 Interatomic potentials and atomistic simulations
An interatomic potential $V$ describes the potential energy between the $N$ atoms of a molecule or material as a function of their positions $\{\mathbf{r}_{i}\}_{i=1}^{N}$ and chemical nature $\{a_{i}\}_{i=1}^{N}$:
@@
This technique consists of finding one or more minima of the potential energy surface $V$ to identify the preferred structure of a microscopic system. Both local and global optimization algorithms are employed, and the vast majority use the forces $\mathbf{f}_{j}$ both during optimization and as a stopping criterion.
@@
MD in the NVE ensemble. The number of atoms $N$, volume $V$ and total energy $E$ (kinetic plus potential) are kept constant, describing the behavior of an isolated system. Usually, this is accomplished by using the velocity Verlet integrator (Verlet, 1967) with a short time step $\Delta t$.
@@ 2.2 Physical symmetries
Interatomic potentials are invariant under the transformations of the Euclidean group in three dimensions E(3), including translations, rotations, and reflections. Mathematically, given a group element $g\in E(3)$ that acts on positions $\mathbf{r}$,
@@
Interatomic potentials are invariant with respect to permutations of atoms, i.e.,
@@ 2.4 Breaking physical constraints
The conservative property for forces in Eq. (5) is satisfied if and only if there exists a function (usually named the potential energy) of which the forces are the spatial derivatives, see Eq. (2). Therefore, a machine learning model that calculates interatomic forces as derivatives of the potential energy according to Eq. (2) will be trivially conservative. The required derivatives, the gradient of the scalar $V$, can be obtained efficiently (i.e., with the same asymptotic computational cost as the energy prediction) with automatic differentiation (Griewank & Walther, 2008). Nevertheless, differentiation incurs a computational overhead, typically $2\mathrm{-}3\times$ for inference and $3\times$ for training; the exact theoretical factors are discussed further in Appendix B. This overhead can be avoided by directly predicting forces during the forward pass. However, by removing the relationship between forces and energy defined in Eq. (2), such models do not enforce energy conservation. The possibility of performing a direct, non-conservative, evaluation of the forces was realized early (Li et al., 2015), but used only sparsely until recently, when it was applied to equivariant graph neural networks such as GemNet (Gasteiger et al., 2021) and several more recent universal machine learning interatomic potentials including ORB (Neumann et al., 2024) and Equiformer (Liao et al., 2023), which are trained on large datasets and promise broad applicability for chemical modeling.
@@ 2.5 Use of forces for training
One important consideration is that focusing on energies or forces emphasizes different components of a potential energy surface, with consequences that are not directly visible when inspecting training and validation accuracies. To a first approximation, the distortions observed in NVT MD simulations – a common strategy to build training sets – can be interpreted as a collection of quasi-harmonic oscillators of different frequency $\omega$, with high-frequency modes associated with short-range covalent bonding, and low-frequency ones associated with collective motions, which are often the most relevant for applications. The statistical mechanics of a harmonic oscillator imply that $\langle V\rangle\propto 1$, while $\langle f^{2}\rangle\propto\omega^{2}$ (see Appendix A). In other terms, the largest contributions to the forces from a thermally sampled dataset come from high-frequency modes that are hard to integrate and may lead to instabilities in the dynamics, while the contribution from the slow modes, which are hard to sample, but usually the most relevant, are under-emphasized. The potential energy instead provides a more balanced representation of the different molecular time and length scales.
@@ 3.1 Measuring non-conservative behavior
where $\mathbf{J}_{\mathrm{anti}}=(\mathbf{J}-\mathbf{J}^{\top})/2$. $\lambda$ then defines a metric going from 0 for conservative forces to $1$ for forces that have no conservative component. $\lambda$ can also be computed only for entries associated with an atom $i$, or a pair of atoms $ij$, providing a finer-grained assessment of the violation of Eq. 6, as seen in Figure 1.
@@ 3.3 Effects on geometry optimization
In order to compare the stability of materials or molecules at low temperature, a common approximation is to search for minimum-energy configurations. This can be achieved by minimizing the potential energy $V$ as a function of the atomic positions – with most of the widely used algorithms relying (in some cases exclusively) on the value of the gradient. The lack of a consistent potential energy is problematic for most optimization schemes: those which require an explicit evaluation of the objective function (e.g., to perform line searches) cannot be used; those that just “follow the forces”, relying on the vanishing magnitude of the force as a stopping criterion, can fail because non-conservative forces can keep driving indefinitely in the same direction, e.g., following closed loops with negative total work.
@@ 3.4 Effects on molecular dynamics
It is not uncommon to observe violations of energy conservation in MD simulations, since finite-timestep integrators violate the exact correspondence between the change in kinetic and potential energy along a trajectory. This is usually accepted, because (1) in well-designed simulations this leads to small fluctuations, and not to a run-away effect; (2) the notion of a shadow Hamiltonian (Hairer et al., 2003) ensures that simulations reach a steady state that is “statistically close” to that generated by an exact integrator; (3) thermostatting techniques can relatively easily control small integration errors, so that structural and dynamical observables are not affected significantly (Morrone et al., 2011). The fact that no underlying Hamiltonian can be defined for the dynamics generated by a non-conservative force field suggests that, in this case, artifacts might be more pronounced and harder to correct. As we shall see, this is what we observe empirically.
@@ 4.1 The models
In order to substantiate our empirical observations, we perform our experiments on multiple models. Our main examples rely on the rotationally unconstrained PET architecture (Pozdnyakov & Ceriotti, 2023), trained from scratch on the bulk water dataset of Ref. 17 using both a conservative (PET) and non-conservative (PET-NC) architecture. Additionally, we train “PET-M” to predict both conservative and non-conservative forces. To assess the implications of a direct prediction of forces in the most favorable possible context, we primarily use the best-performing, custom-trained models. We also show some results for the non-conservative ORB model (Neumann et al., 2024), which is currently state-of-the-art for several materials prediction benchmarks. Even though ORB is not trained on this specific dataset, and is therefore at a clear disadvantage, it provides an indication of the relevance of the issues we discuss. In the appendices, we also discuss several other architectures, including a “legacy” SOAP-BPNN architecture, as well as pre-trained foundation models, MACE-MP-0 (Batatia et al., 2023), SevenNet (Park et al., 2024), and EquiformerV2 (Liao et al., 2023). A table describing all employed models, along with full details on the different architectures, can be found in Appendix C.
Table 1: Test errors (energies in meV per atom, forces in meV/Å) of PET models trained on the bulk water dataset.
@@ 4.2 Accuracy
In terms of sheer accuracy, our tests confirm that forces provide very useful information to train an interatomic potential, in particular for a dataset containing relatively large configurations. Using forces in the training of a conservative model dramatically improves the accuracy of energy predictions with just a minor degradation in the accuracy for $\mathbf{f}$ with respect to a model trained only on forces. A non-conservative force model exhibits about 30% higher force error than a conservative architecture, and including a separate energy head leads to lower error than a model trained just on $V$ – indicating that the sharing of weights within the architecture is beneficial. We also show results for the PET-M hybrid architecture, which makes both conservative and non-conservative force predictions. Its accuracy is less than 10% worse than the best models for either architectures. As we will discuss in Section 4.8, this is an excellent way to exploit non-conservative forces in simulations. Before doing so, however, we will assess the behavior of purely non-conservative models.
@@ 4.3 Non-conservative behavior
The asymmetry of the Jacobian is the most direct, pointwise measure of non-conservative behavior. Different non-conservative models show widely different values of $\lambda$ – 0.015 for ORB, 0.017 for Equiformer, 0.032 for SOAP-BPNN-NC and 0.004 for PET-NC, computed on a few water structures from the test set of Cheng et al. (2019). As we shall see, the magnitude of $\lambda$ correlates qualitatively with the stability of the models in simulations. The symmetry of $\mathbf{J}$ applies separately to each pair of atoms, and so it is possible to extract further insights by computing the norm of the antisymmetric part of each block $\mathbf{J}_{ij}$ resolved for different atomic pairs and plotted as a function of the interatomic distance (Figure 1). The asymmetry is also present for “on-site” blocks, i.e., swapping only the Cartesian coordinates used in the derivatives for a given atom; the relative magnitude of $\mathbf{J}_{\text{asym}}$ is small compared to the magnitude of the Jacobian; the decay of the asymmetric component between atoms $i$ and $j$ with the interatomic distance decays more slowly than the absolute magnitude of $\mathbf{J}$ (which has been used as a measure of the interactions between pairs of atoms (Herbold & Behler, 2022)), and in the intermediate regime around 6Å it becomes comparable in size – so that the pair-resolved Jacobian asymmetry $\lambda_{ij}$ approaches 1 for large interatomic distances. This latter observation has important implications when applying these models in simulations, as the impact of non-conservative behavior on different atomic-scale processes is not uniform, and it tends to be larger – in a relative sense – for collective processes involving long-range correlations.
@@
Non-conservative behavior can also be demonstrated by computing the work along a closed path (see Appendix D). Given that the choice of the path is arbitrary, we think it is more relevant to quantify the practical implications of this issue in terms of an energy drift in molecular simulations.
@@ 4.4 Constant-energy molecular dynamics
Let us now consider the use of non-conservative MLIPs in the context of MD simulations. Also in this case, we focus our analysis on the best-performing models for liquid water, PET and PET-NC, leaving a more thorough comparison that consistently corroborates the observations we make here for Appendix F. Given that non-conservative models lack a well-defined conserved quantity by construction, we rely on indirect measurements of the sampled ensemble. We consider in particular the kinetic temperature $T=2K/(3Nk_{B})$, where $N$ is the number of particles considered. This is just a rescaling of the kinetic energy $K$; its ensemble average should correspond to the target temperature for NVT trajectories (300 K in these tests), and to a value in its vicinity for NVE trajectories initialized from a thermally equilibrated configuration. As a sensitive indicator of the dynamical behavior of the system, we compute the Fourier transform of the velocity-velocity correlation function, $\hat{c}_{vv}(\omega)$. Its peaks are closely related to the density of vibrational modes and to infrared and Raman spectra, and its $\omega\rightarrow 0$ limit is proportional to the diffusion coefficient.
@@
The failure of the non-conservative model in NVE dynamics is apparent in Figure 2. Whereas for a conservative potential the kinetic temperature fluctuates around the initial value, the spurious work associated with non-conservative forces leads to a large drift of $T$: To put it on a human scale, this unphysical drift corresponds to a rate of heating of about 7’000 billion degrees per second for the custom-trained PET-NC model, and another 10 times larger for the general-purpose ORB model. This spurious energy flow is a clear signature of non-conservative behavior, and makes direct-force models entirely useless for constant-energy simulations.

### 27 May 2025 v3

@@ Abstract
The use of machine learning to estimate the energy of a group of atoms, and the forces that drive them to more stable configurations, has revolutionized the fields of computational chemistry and materials discovery. In this domain, rigorous enforcement of symmetry and conservation laws has traditionally been considered essential. For this reason, interatomic forces are usually computed as the derivatives of the potential energy, ensuring energy conservation. Several recent works have questioned this physically constrained approach, suggesting that directly predicting the forces yields a better trade-off between accuracy and computational efficiency – and that energy conservation can be learned during training. This work investigates the applicability of such non-conservative models in microscopic simulations. We identify and demonstrate several fundamental issues, from ill-defined convergence of geometry optimization to instability in various types of molecular dynamics. Contrary to the case of rotational symmetry, energy conservation is hard to learn, monitor, and correct for. The best approach to exploit the acceleration afforded by direct force prediction might be to use it in tandem with a conservative model, reducing – rather than eliminating – the additional cost of backpropagation, but avoiding the pathological behavior associated with non-conservative forces.
Machine Learning, ICML, Molecular Dynamics, Interatomic Potentials, Statistical Mechanics, Geometric Machine Learning
@@ 1 Introduction
Interatomic potentials model the microscopic interactions between atoms, and determine – directly or by modulating the response to thermal excitations – the stability and reactivity of molecules and materials. Over many decades, interatomic potentials have been used in Monte Carlo (Metropolis et al., 1953) and molecular dynamics (MD) (Andersen, 1980) simulations, geometry optimization, and other techniques, allowing a mechanistic study of the atomic-scale behavior and properties of molecules, materials, and biological systems (Allen & Tildesley, 2017; Tuckerman, 2008). While traditional interatomic potentials are based on simple physically inspired functional forms, the last decade has seen machine-learned interatomic potentials (MLIPs) obtain remarkable accuracies at a high level of computational efficiency, most often by learning reference energies and forces from much slower first-principle quantum mechanical calculations (Behler, 2021; Unke et al., 2021).
@@ 2.1 Interatomic potentials and atomistic simulations
An interatomic potential $V$ describes the potential energy between the NN atoms of a molecule or material as a function of their positions $\{\mathbf{r}_{i}\}_{i=1}^{N}$ and chemical nature $\{a_{i}\}_{i=1}^{N}$:
@@
This technique consists of finding one or more minima of the potential energy surface $V$ to identify the preferred structure of a microscopic system. Both local and global optimization algorithms are employed, and the vast majority use the forces $\mathbf{f}_{j}$ both during optimization and as a stopping criterion. Similar methods are used to identify saddle points of $V$, which are associated with the energy barriers that determine the rate of chemical reactions (Henkelman et al., 2000).
@@
MD in the NVE ensemble. The number of atoms NN, volume $V$ and total energy $E$ (kinetic plus potential) are kept constant, describing the behavior of an isolated system. Usually, this is accomplished by using the velocity Verlet integrator (Verlet, 1967) with a short time step $\Delta t$.
@@ 2.2 Physical symmetries
Interatomic potentials are invariant under the transformations of the Euclidean group in three dimensions E(3), including translations, rotations, and reflections. Mathematically, given a group element $g\in E(3)$ that acts on all positions $\mathbf{r}$,
@@
Interatomic potentials are invariant with respect to permutations of atom indices, i.e.,
@@ 2.4 Breaking physical constraints
The conservative property for forces in Eq. (5) is satisfied if and only if there exists a function (usually named the potential energy) of which the forces are the spatial derivatives, see Eq. (2). Therefore, a machine learning model that calculates interatomic forces as derivatives of the potential energy according to Eq. (2) will be trivially conservative. The required derivatives, the gradient of the scalar $V$, can be obtained efficiently (i.e., with the same asymptotic computational cost as the energy prediction) with automatic differentiation (Griewank & Walther, 2008). Nevertheless, differentiation incurs a computational overhead, typically $2\mathrm{-}3\times$ for inference and $3\times$ for training; the exact theoretical factors are discussed further in Appendix B. This overhead can be avoided by directly predicting forces during the forward pass. However, by removing the relationship between forces and energy defined in Eq. (2), such models do not enforce energy conservation. The possibility of performing a direct, non-conservative, evaluation of the forces was realized early (Li et al., 2015), but used only sparsely until recently, when it was applied to equivariant graph neural networks such as GemNet (Gasteiger et al., 2021) and several more recent universal machine learning interatomic potentials including ORB (Neumann et al., 2024) and Equiformer (Liao et al., 2023), which are trained on large datasets and promise broad applicability for chemical modeling. Some of these architectures also perform direct evaluation of stresses, which leads to the violation of conservation of enthalpy (see App. G.3).
The impact of lack of energy conservation for practical simulations and property prediction has recently been investigated in related works: Eissler et al. (2025) probe the limits of unconstrained architectures and find that lack of energy conservation becomes more pronounced for larger target systems. Loew et al. (2024) find that direct-force models perform poorly at the prediction of phonon properties, which require second derivatives of the potential energy surface. Fu et al. (2025) observe a better correlation between test set error and downstream predictive performance, for instance for phonons, for models that can perform stable MD simulations, i.e., that are able to conserve energy.
@@ 2.5 Use of forces for training
One important consideration is that focusing on energies or forces emphasizes different components of a potential energy surface, with consequences that are not directly visible when inspecting training and validation accuracies. To a first approximation, the distortions observed in NVT MD simulations – a common strategy to build training sets – can be interpreted as a collection of quasi-harmonic oscillators of different frequency $\omega$, with high-frequency modes associated with short-range covalent bonding, and low-frequency ones associated with collective motions, which are often the most relevant for applications. The statistical mechanics of a harmonic oscillator imply that $\langle V\rangle\propto 1$, while $\langle f^{2}\rangle\propto\omega^{2}$ (see Appendix A). In other terms, the largest contributions to the forces from a thermally sampled dataset come from high-frequency modes that are hard to integrate and may lead to instabilities in the dynamics, while the contribution from the slow modes, which are hard to sample, but usually the most relevant, are under-emphasized. The potential energy provides a more balanced representation of the different molecular time and length scales.
@@ 3.1 Measuring non-conservative behavior
where $\mathbf{J}_{\mathrm{anti}}=(\mathbf{J}-\mathbf{J}^{\top})/2$. $\lambda$ then defines a metric going from 0 for conservative forces to 11 for forces that have no conservative component. $\lambda$ can also be computed only for entries associated with an atom $i$, or a pair of atoms ijij, providing a finer-grained assessment of the violation of Eq. 6, as seen in Figure 1.
@@ 3.2 Side-effects of direct force prediction
There is another non-trivial consequence of using a direct-force prediction architecture. Whereas the potential energy is usually estimated as the sum of atomic contributions but is a global property, forces are atom-centered. When predicting them as derivatives, many atomic environments contribute to the force $\mathbf{f}_{i}$ on each atom. When predicting directly, only the $i$-atom centered environment contributes to $\mathbf{f}_{i}$. Hence, direct force models can be expected to be affected more directly by the geometric degeneracies of low-body-order atom-centered descriptors (Pozdnyakov et al., 2020) and do not benefit from the same extended effective interaction range as conservative forces (Artrith et al., 2011). We discuss these effects in Appendix C, presenting some empirical evidence that direct force models require a larger range to match the force accuracy of a comparable conservative model.
@@ 3.3 Effects on geometry optimization
In order to assess the stability of materials or molecules at low temperature, a common approximation is to search for minimum-energy configurations. This can be achieved by minimizing the potential energy $V$ as a function of the atomic positions – with most of the widely used algorithms relying (in some cases exclusively) on the value of the gradient. The lack of a consistent potential energy is problematic for most optimization schemes: those which require an explicit evaluation of the objective function (e.g., to perform line searches) cannot be used; those that just “follow the forces”, relying on the vanishing magnitude of the force as a stopping criterion, can fail because non-conservative forces can keep driving indefinitely in the same direction, e.g., following closed loops with negative total work.
@@ 3.4 Effects on molecular dynamics
It is not uncommon to observe violations of energy conservation in MD simulations, since finite-timestep integrators violate the exact correspondence between the change in kinetic and potential energy along a trajectory. This is usually accepted, because (1) in well-designed simulations this leads to small fluctuations, and not to a run-away effect; (2) the notion of a shadow Hamiltonian (Hairer et al., 2006) ensures that simulations reach a steady state that is “statistically close” to that generated by an exact integrator; (3) thermostatting techniques can relatively easily control small integration errors, so that structural and dynamical observables are not affected significantly (Morrone et al., 2011). The fact that no underlying Hamiltonian can be defined for the dynamics generated by a non-conservative force field suggests that, in this case, artifacts might be more pronounced and harder to correct. For example, the symplectic nature of Hamiltonian dynamics is no longer valid (this is easy to see, for example, from Eq. 5.2, Chapter 6 of Hairer et al. (2006)), and the theorem of equipartition of energy (whose proof is also based on the existence of a Hamiltonian (Pathria, 2017)) does not apply. As we shall see, this is what we observe empirically.
@@ 4.1 The models
In order to substantiate our empirical observations, we perform our experiments on multiple models. Our main examples rely on the rotationally unconstrained PET architecture (Pozdnyakov & Ceriotti, 2023), trained from scratch on the bulk water dataset of Ref. 22 using both a conservative (PET) and non-conservative (PET-NC) architecture. Additionally, we train “PET-M” to predict both conservative and non-conservative forces. To assess the implications of a direct prediction of forces in the most favorable possible context, we primarily use the best-performing, custom-trained models. We also show some results for the non-conservative ORB-v2 model (Neumann et al., 2024), which is currently state-of-the-art for several materials prediction benchmarks. Even though ORB is not trained on this specific dataset, and is therefore at a clear disadvantage, it provides an indication of the relevance of the issues we discuss. In the appendices, we also discuss several other architectures, including a “legacy” SOAP-BPNN architecture, as well as pre-trained foundation models, MACE-MP-0 (Batatia et al., 2023), SevenNet (Park et al., 2024), and EquiformerV2 (Liao et al., 2023), which we apply to a few diverse materials in Appendix G. A table describing all employed models, along with full details on the different architectures, can be found in Appendix D.
Table 1: Test errors (energies in meV per atom, forces in meV/Å) of PET models trained on the bulk water dataset. In this table, and all subsequent tables, rows corresponding to non-conservative models are highlighted with a gray background.
@@ 4.2 Accuracy
In terms of sheer accuracy, see Table 1, our tests confirm that forces provide very useful information to train an interatomic potential, in particular for a dataset containing relatively large configurations. Using forces in the training of a conservative model dramatically improves the accuracy of energy predictions with just a minor degradation in the accuracy for $\mathbf{f}$ with respect to a model trained only on forces. A non-conservative force model exhibits about 30% higher force error than a conservative architecture, and including a separate energy head leads to lower error than a model trained just on $V$ – indicating that the sharing of weights within the architecture is beneficial. We also show results for the PET-M hybrid architecture, which makes both conservative and non-conservative force predictions. Its accuracy is less than 10% worse than the best models for either architectures. As we will discuss in Section 4.8, this is an excellent way to exploit non-conservative forces in simulations. Before doing so, however, we will assess the behavior of purely non-conservative models.
@@ 4.3 Non-conservative behavior
The asymmetry of the Jacobian is the most direct, pointwise measure of non-conservative behavior. Different non-conservative models show widely different values of $\lambda$ – 0.015 for ORB, 0.017 for Equiformer, 0.032 for SOAP-BPNN-NC and 0.004 for PET-NC, computed on a few water structures from the test set of Cheng et al. (2019). As we shall see, the magnitude of $\lambda$ correlates qualitatively with the stability of the models in simulations. The symmetry of $\mathbf{J}$ applies separately to each pair of atoms, and so it is possible to extract further insights by computing the norm of the antisymmetric part of each block $\mathbf{J}_{ij}$ resolved for different atomic pairs and plotted as a function of the interatomic distance (Figure 1). The asymmetry is also present for “on-site” blocks, i.e., swapping only the Cartesian coordinates used in the derivatives for a given atom; the relative magnitude of $\mathbf{J}_{\text{asym}}$ is small compared to the magnitude of the Jacobian; the asymmetric component between atoms $i$ and $j$ decays with the interatomic distance more slowly than the absolute magnitude of $\mathbf{J}$ (which has been used as a measure of the interactions between pairs of atoms (Herbold & Behler, 2022)), and in the intermediate regime around 6Å it becomes comparable in size – so that the pair-resolved Jacobian asymmetry $\lambda_{ij}$ approaches 1 for large interatomic distances. This latter observation has important implications when applying these models in simulations, as the impact of non-conservative behavior on different atomic-scale processes is not uniform, and it tends to be larger – in a relative sense – for collective processes involving long-range correlations.
@@
Non-conservative behavior can also be demonstrated by computing the work along a closed path (see Appendix E). Given that the choice of the path is arbitrary, we think it is more relevant to quantify the practical implications of this issue in terms of an energy drift in molecular simulations.
@@ 4.4 Constant-energy molecular dynamics
Let us now consider the use of non-conservative MLIPs in the context of MD simulations. Also in this case, we focus our analysis on the best-performing models for liquid water, PET and PET-NC. A more thorough comparison, including several foundation models and a few homogeneous and heterogeneous material structures, is discussed in Appendix G, and consistently corroborates the observations we make here. Constant-pressure simulations, which, if performed with direct-stress models, break conservation of enthalpy and therefore lead to drift in the volume of the simulation, are shown in App. G.3 using the general-purpose PET-MAD potential (Mazitov et al., 2025).
Given that non-conservative models lack a well-defined conserved quantity by construction, we rely on indirect measurements of the sampled ensemble. We consider in particular the kinetic temperature $T=2K/(3Nk_{\mathrm{B}})$, where NN is the number of particles considered. This is just a rescaling of the kinetic energy $K$; its ensemble average should correspond to the target temperature for NVT trajectories (300 K in these tests), and to a value in its vicinity for NVE trajectories initialized from a thermally equilibrated configuration. As a sensitive indicator of the dynamical behavior of the system, we compute the Fourier transform of the velocity-velocity correlation function, $\hat{c}_{vv}(\omega)$. Its peaks are closely related to the density of vibrational modes and to infrared and Raman spectra, and its $\omega\rightarrow 0$ limit is proportional to the diffusion coefficient.
@@
The failure of the non-conservative model in NVE dynamics is apparent in Figure 2. Whereas for a conservative potential the kinetic temperature fluctuates around the initial value, the spurious work associated with non-conservative forces leads to a large drift of $T$: To put it on a human scale, this unphysical drift corresponds to a rate of heating of about 7’000 billion degrees per second for the custom-trained PET-NC model, and another 10 times larger for the general-purpose ORB model. This spurious energy flow is a clear signature of non-conservative behavior, and makes direct-force models entirely useless for constant-energy simulations. This runaway increase of the kinetic energy can be mitigated – whenever an auxiliary model is available to evaluate the potential energy – by adjusting the velocities to enforce energy conservation artificially (see Appendix G.4). Similar to what we will discuss for constant-temperature simulations, the trajectories are still affected by large artefacts.